# Unveiling Impact of Frequency Components on Membership Inference Attacks for Diffusion Models

### v2

@@ 1 Introduction
Diffusion models (Song et al., 2020a; Ho et al., 2020) have achieved significant success in areas such as image generation (Choi et al., 2021; Si et al., 2024; Huberman-Spiegelglas et al., 2024) and video generation (Wu et al., 2024; Blattmann et al., 2023), and have been widely applied in real life. However, this success has brought increasing attention to copyright and privacy issues from both academia and industry (Wang et al., 2023; Wen et al., 2024; Zhang et al., 2024). Recent research shows that diffusion models exhibit a strong memory effect regarding images in the training set, making the risk of privacy leakage a serious concern (Wen et al., 2024; Duan et al., 2023).
@@
Recently, MIAs for diffusion models have gradually emerged as a research hotspot. Matsumoto et al. (2023) proposed a query-based attack strategy that determines member data by analyzing the model’s loss function. SecMI (Duan et al., 2023) is based on DDIM inversion (Song et al., 2020a; Kim et al., 2022), obtaining intermediate outputs during generation through a deterministic inversion process. Kong et al. (2023) introduced a proximal initialization method that utilizes model predictions to obtain initial noise. Zhai et al. (2024) studied text-to-image diffusion models, exploring the associative memory between texts and images based on likelihood estimation. In summary, the core mechanism of these methods is to quantify the model’s image recovery ability and use this as the basis for constructing the membership inference decision logic.
@@
  • To the best of our knowledge, this study is the first to explore the impact of frequency domain information on MIAs targeting diffusion models. We formalize a general paradigm for existing error-based attacks and conduct an in-depth analysis of the impact of frequency domain information. The results reveal that existing attacks generally overlook the standard deviation induced by high-frequency deficiency, which restricts their attack performance.
@@ 2 Related Works
Membership Inference Attacks. Shokri et al. (2017) proposed the membership inference attacks, which primarily targeted classification models in machine learning. As the evolution of membership inference attacks continues, they can be classified into two categories: black-box attacks (Salem et al., 2018; Song & Mittal, 2021; Choquette-Choo et al., 2021) and white-box attacks (Nasr et al., 2019; Leino & Fredrikson, 2020), determined by the degree of access granted to the target model. In a white-box setting, the attacker can access the model parameters, whereas in a black-box setting, the attacker only receives the final output of the model. Moreover, noteworthy advancements have been achieved in membership inference attacks targeting generative models. Hayes et al. (2017) demonstrated that membership can be effectively discerned through the logits of the discriminator in GANs. Hilprecht et al. (2019) introduced a Monte Carlo scoring methodology incorporating the reconstruction loss term to facilitate attacks on VAEs.
Membership Inference Attacks on Diffusion Models. Recently, membership inference attacks on diffusion models have garnered increasing attention. In white-box settings, Pang et al. (2023) proposed executing an attack through the utilization of gradient information extracted from loss. In grey-box settings, the attacker can only access the intermediate and final outputs (Duan et al., 2023). Matsumoto et al. (2023) pioneered the approach of employing diffusion loss to perform query-based membership inference. Duan et al. (2023) introduced an attack leveraging DDIM inversion to retrieve intermediate outputs from the models. Meanwhile, Kong et al. (2023) proposed a proximal initialization technique to acquire the deterministic initial noise. The attack is realized by the prediction of this noise from the model. In addition, attacks leveraging the correlation between texts and images have also achieved advancements (Wen et al., 2024; Zhai et al., 2024; Li et al., 2024). Moreover, there has also been a growing interest in attacks for diffusion models in black-box settings (Pang & Wang, 2023).
@@ 3 Preliminaries
To understand how frequency information affects diffusion models’ behaviour in membership inference contexts, we first establish the fundamental mechanisms of diffusion models and how images can be represented in the frequency domain.
@@
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}\left(\sqrt{\overline{\alpha}_{t-1}}x_{0}+\sqrt{1-\overline{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\frac{x_{t}-\sqrt{\overline{\alpha}_{t}}x_{0}}{\sqrt{1-\overline{\alpha}_{t}}},\sigma_{t}^{2}I\right),$ (1)
@@
$x_{t-1}=\sqrt{\overline{\alpha}_{t-1}}\left(\frac{x_{t}-\sqrt{1-\overline{\alpha}_{t}}\epsilon_{\theta}(x_{t},t)}{\sqrt{\overline{\alpha}_{t}}}\right)+\sqrt{1-\overline{\alpha}_{t-1}-\sigma_{t}^{2}}\epsilon_{\theta}(x_{t},t)+\sigma_{t}\epsilon,$ (2)
where $\epsilon\sim\mathcal{N}(0,I)$, $\sigma_{t}=\eta\sqrt{(1-\overline{\alpha}_{t-1})/(1-\overline{\alpha}_{t})}\sqrt{1-\overline{\alpha}_{t}/\overline{\alpha}_{t-1}}$, $\eta\in[0,1]$. The case $\eta=0$ corresponds to the deterministic DDIM, while $\eta=1$ corresponds to the DDPM. More details about the DDPM is provided in Appendix A.1.
@@
$\mathbf{X}(u,v)=FFT(\mathbf{x})=\sum_{x=1}^{H}\sum_{y=1}^{W}\mathbf{x}(x,y)e^{-j2\pi\left(\frac{u}{H}x+\frac{v}{W}y\right)},$ (3)
@@
$\mathbf{x}(x,y)=IFFT(\mathbf{x})=\frac{1}{HW}\sum_{u=1}^{H}\sum_{v=1}^{W}\mathbf{X}(u,v)e^{j2\pi\left(\frac{u}{H}x+\frac{v}{W}y\right)}.$ (4)
@@ 4.1 Formalization of MIAs for Diffusion Models
General Paradigm. Attacks such as Naive (Matsumoto et al., 2023), SecMI (Duan et al., 2023) and PIA (Kong et al., 2023), relying on reconstruction errors, have demonstrated effectiveness in diffusion models. These attacks share common characteristics, utilizing the model’s image prediction capability to execute attacks. They can be unified as follows: given the image to be tested $x_{i}$, calculate the distance between the image $x_{i,t}$ predicted by the model at step $t$ and the target result $x_{i,t}^{target}$ of $x_{i}$ at step $t$, then using the distance as the membership score. $x_{i,t}$ and $x_{i,t}^{target}$ are obtained in different ways depending on the algorithm. Furthermore, setting a threshold $\tau$. If the score is less than $\tau$, the image is classified as member data; otherwise, it is classified as hold-out data. Formally, these attacks can be formulated as the general paradigm:
@@
4.2 Frequency Perspective of MIAs for Diffusion Models
Frequency Characteristics of Diffusion Models. Existing attack studies on diffusion models mainly concentrate on reconstruction errors, yet they neglect an important aspect: analyzing models’ information processing from the frequency domain. Diffusion models’ operational mechanism features distinct frequency hierarchical properties. They first denoise low-frequency signals according to the learned distribution, then utilize specific low-frequency information as prior knowledge to process high-frequency details (Qian et al., 2024). Recovering high-frequency information effectively relies not only on the model’s learned distribution but also closely ties to the image’s inherent structure and contours during denoising. Previous research (Yang et al., 2023) has shown that diffusion models exhibit more variation and uncertainty in handling high-frequency information.
@@ 4.2 Frequency Perspective of MIAs for Diffusion Models
Frequency Effects on MIAs for Diffusion. To study how frequency domain information affects MIAs, we visualized the relationship between high-frequency content and existing attacks’ membership scores. First, we transform images to the frequency domain using the Fourier transform and divide the high- and low-frequency regions by setting the frequency domain radius =5=5 as the high-low frequency boundary. Then, we calculated the percentage of high-frequency content by summing the squared frequency components in both regions. As shown in Fig. 1, visualizing the scores of attacks reveals a trend that as the high-frequency content of the images increased, so did the membership scores. This shows current attacks are biased, giving higher scores to images with more high-frequency content. Higher scores mean a lower degree of fitting, making it more likely to be classified as hold-out data. We term the phenomenon "high-frequency deficiency".
Table 1: High-frequency content statistics for failed samples. In the failed samples, the high-frequency content of the member data is significantly larger than that of the hold-out data.
Method
STLU-10 Tiny-IN MS-COCO Flickr
@@
4.3 High-Frequency Filter Design for Enhanced Diffusion MIAs
Based on the above observations, the following conclusion can be drawn: existing MIAs have not adequately accounted for the effect of high-frequency deficiency, which leads to their failure in certain specific scenarios. Intuitively, the inherent deficiency of diffusion models in handling high-frequency components leads to confusion in distinguishing between images with different high-frequency content. Further, we explore the effects of high-frequency deficiency from a theoretical perspective. We rely on the previously established definition of membership inference attack capability (Yeom et al., 2018). Definition 1 gives an advantage measure that characterizes how well an algorithm can distinguish between member data and hold-out data.
Definition 1.
The membership advantage of algorithm $\mathcal{A}$ is defined as:
@@ 4.3 High-Frequency Filter Design for Enhanced Diffusion MIAs
where $\Pr[\mathcal{A}=1|m=1]$ indicates the probability that algorithm $\mathcal{A}$ identifies the data as a member when the data is a member. $\Pr[\mathcal{A}=1|m=0]$ indicates the probability that if the data is hold-out data, algorithm $\mathcal{A}$ will identify it as a member.
When membership scores follow a normal distribution, Yeom et al. (2018) reinterpret the membership advantage in terms of distributional differences. Specifically, they express the advantage using the ratio between the standard deviation of the hold-out scores $\sigma_{H}$ and that of the member scores $\sigma_{M}$. And they further demonstrate that this ratio is positively correlated with the membership advantage:
@@
A larger ratio $\sigma_{H}/\sigma_{M}$ corresponds to a greater advantage for the attacker. When $\sigma_{H}<\sigma_{M}$, it means that there is no member advantage. Mathematically, the high-frequency deficiency contributes equally to a high standard deviation in the scores of both member and hold-out data, which may lead to a decrease in $\sigma_{H}/\sigma_{M}$, weakening the advantages of identifying member data and consequently interfering with the effective differentiation between member and hold-out data. This perspective is further validated in proposition 1.
@@
$\beta_{i,t}(r)=\begin{cases}s&\text{if }r>r_{t},\\ 1&\text{otherwise.}\end{cases}$ (10)
@@
Proposition 1.
Denote the original standard deviations of membership scores in member and hold-out data as $\sigma_{M}$ and $\sigma_{H}$, and the standard deviations after removing the high-frequency components are $\sigma_{M}^{\prime}$ and $\sigma_{H}^{\prime}$. The standard deviation of membership scores in the high-frequency components is $h_{M}(h_{H})$, and the low-frequency components is $l_{M}(l_{H})$ in member(hold-out) data. Let $l_{H}-l_{M}=\Delta$, $h_{M}=k\cdot h_{H}$ with $k>0$. If $k^{2}>1+\frac{2\Delta}{h_{H}^{2}}(l_{M}+2\Delta-\sqrt{(l_{M}+2\Delta)^{2}+h_{H}^{2}})$, we have:
@@
We provide the detailed proof of Proposition 1 in Appendix C. ∎
Since $l_{M}+2\Delta-\sqrt{(l_{M}+2\Delta)^{2}+h_{H}^{2}}<0$, $\sigma_{H}^{\prime}/\sigma_{M}^{\prime}>\sigma_{H}/\sigma_{M}$ is hold constantly when $k\geq 1$. The key insight of Proposition 1 is that when the ratio of $h_{M}/h_{H}$ exceeds a certain threshold, high-frequency contents will weaken the membership advantage, and filtering the membership scores derived from high-frequency components will amplify the membership advantage. From a theoretical perspective, it demonstrates that high-frequency components reduce the attack capability of the algorithm, while also validating the effectiveness of high-frequency filtering methods.
To further validate the practical applicability of the theoretical framework, we investigated the constraint conditions of $k$. The result shows that $k$ satisfies its constraint conditions under normal circumstances. We also provide a discussion on normality and verify the validity of the normality assumption in Proposition 1, while experimental results further demonstrate that our method does not rely on the assumption of normality. Detailed experiments and analysis are provided in Appendix D.2.
5 Experiments
5.1 Experiment Setup
Baselines. We employed the SecMI (Duan et al., 2023), PIA (Kong et al., 2023), and Naive (Matsumoto et al., 2023) as our baselines for comparison. These approaches are characterized as error-based attacks and operate effectively within a grey-box setting. We adopt the parameters advised in their respective publications.
@@ 5.1 Experiment Setup
Method
STL10-U CIFAR-100 Tiny-IN
@@
Method
Pokémon MS-COCO Flickr

### v3

# Enhancing Membership Inference Attacks on Diffusion Models from a Frequency-Domain Perspective

@@ 1 Introduction
Diffusion models (Song et al., 2020a; Ho et al., 2020) have achieved significant success in areas such as image generation (Choi et al., 2021; Si et al., 2024; Huberman-Spiegelglas et al., 2024) and video generation (Wu et al., 2024; Blattmann et al., 2023), and have been widely applied in real life. However, this success has brought increasing attention to copyright and privacy issues from both academia and industry (Wang et al., 2023; Wen et al., 2024; Zhang et al., 2024a). Recent research shows that diffusion models exhibit a strong memory effect regarding images in the training set, making the risk of privacy leakage a serious concern (Wen et al., 2024; Duan et al., 2023).
@@
Recently, MIAs for diffusion models have received increasing attention. (Matsumoto et al., 2023) proposed a query-based attack strategy that determines member data by analyzing the model’s loss function. SecMI (Duan et al., 2023) is based on DDIM inversion (Song et al., 2020a; Kim et al., 2022), obtaining intermediate outputs during generation through a deterministic inversion process. (Kong et al., 2023) introduced a proximal initialization method that utilizes model predictions to obtain initial noise. (Zhai et al., 2024) studied the associative memory between texts and images based on likelihood estimation. In summary, the core mechanism of these methods is to quantify the model’s image recovery ability and use this as the basis for constructing the membership inference decision logic.
@@
  • To the best of our knowledge, this study is the first to explore the impact of frequency domain information on MIAs targeting diffusion models. We formalize a general paradigm for existing error-based attacks and conduct an in-depth analysis of the impact of frequency domain information. The results reveal that existing attacks generally overlook the standard deviation of membership score induced by high-frequency deficiency, which restricts their attack performance.
@@ 2 Related Works
Membership Inference Attacks. (Shokri et al., 2017) proposed the membership inference attacks, which primarily targeted classification models in machine learning. As the evolution of membership inference attacks continues, they can be classified into two categories: black-box attacks (Salem et al., 2018; Song and Mittal, 2021; Choquette-Choo et al., 2021) and white-box attacks (Nasr et al., 2019; Leino and Fredrikson, 2020), determined by the degree of access granted to the target model. In a white-box setting, the attacker can access the model parameters, whereas in a black-box setting, the attacker only receives the final output of the model. Moreover, noteworthy advancements have been achieved in membership inference attacks targeting generative models. (Hayes et al., 2017) demonstrated that membership can be effectively discerned through the logits of the discriminator in GANs. (Hilprecht et al., 2019) introduced a Monte Carlo scoring methodology incorporating the reconstruction loss term to facilitate attacks on VAEs.
Membership Inference Attacks on Diffusion Models. Recently, membership inference attacks on diffusion models have garnered increasing attention. In white-box settings, (Pang et al., 2023) proposed executing an attack through the utilization of gradient information extracted from loss. In grey-box settings, the attacker can only access the intermediate and final outputs (Duan et al., 2023). (Matsumoto et al., 2023) pioneered the approach of employing diffusion loss to perform query-based membership inference. (Duan et al., 2023) introduced an attack leveraging DDIM inversion to retrieve intermediate outputs from the models. Meanwhile, (Kong et al., 2023) proposed a proximal initialization technique to acquire the deterministic initial noise. The attack is realized by the prediction of this noise from the model. In addition, attacks leveraging the correlation between texts and images have also achieved advancements (Wen et al., 2024; Zhai et al., 2024; Li et al., 2024). Moreover, there has also been a growing interest in attacks for diffusion models in black-box settings (Pang and Wang, 2023).
Frequency Analysis for Diffusion Models. (Rissanen et al., 2022) observed the implicit spectral bias of diffusion models, which favors generating low-frequency components before high-frequency components. (Yang et al., 2023) has also demonstrated, from the perspective of the power spectrum of natural images, that diffusion models initially learn to recover low-frequency components, and then, in fewer subsequent time steps, learn to recover high-frequency components. This results in greater uncertainty in the restoration of high-frequency details. (Falck et al., 2025) empirically verifies, from a signal-to-noise ratio perspective, that during training, high-frequency components are more rapidly and early masked compared to low-frequency components. This leads to instability and greater variation in the quality of high-frequency generation. In addition, this inherent low-to-high frequency generation pattern observed in diffusion models is evident in both text-to-image and video generation (Yi et al., 2024; Zhang et al., 2024b).
@@ 3 Preliminaries
To understand how frequency information affects diffusion models’ behaviour in membership inference contexts, we establish the fundamental mechanisms of diffusion models and the representation of images in the frequency domain.
@@
$\begin{split}p_{\theta}(x_{t-1}&|x_{t})=\mathcal{N}\left(x_{t-1};\sqrt{\overline{\alpha}_{t-1}}x_{0}\right.\\ &+\left.\sqrt{1-\overline{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\frac{x_{t}-\sqrt{\overline{\alpha}_{t}}x_{0}}{\sqrt{1-\overline{\alpha}_{t}}},\sigma_{t}^{2}I\right),\end{split}$ (1)
@@
$\begin{split}x_{t-1}=\sqrt{\overline{\alpha}_{t-1}}&\left(\frac{x_{t}-\sqrt{1-\overline{\alpha}_{t}}\epsilon_{\theta}(x_{t},t)}{\sqrt{\overline{\alpha}_{t}}}\right)\\ &+\sqrt{1-\overline{\alpha}_{t-1}-\sigma_{t}^{2}}\epsilon_{\theta}(x_{t},t)+\sigma_{t}\epsilon,\end{split}$ (2)
where $\sigma_{t}=\eta\sqrt{(1-\overline{\alpha}_{t-1})/(1-\overline{\alpha}_{t})}\sqrt{1-\overline{\alpha}_{t}/\overline{\alpha}_{t-1}}$, $\eta\in[0,1]$, $\epsilon\sim\mathcal{N}(0,I)$. The case $\eta=0$ corresponds to the DDIM, while $\eta=1$ corresponds to the DDPM. More details about the DDPM are provided in Appendix A.1.
@@
$FFT(\mathbf{x})=\sum_{x=1}^{H}\sum_{y=1}^{W}\mathbf{x}(x,y)e^{-j2\pi\left(\frac{u}{H}x+\frac{v}{W}y\right)},$ (3)
@@
$IFFT(\mathbf{X})=\frac{1}{HW}\sum_{u=1}^{H}\sum_{v=1}^{W}\mathbf{X}(u,v)e^{j2\pi\left(\frac{u}{H}x+\frac{v}{W}y\right)}.$ (4)
@@ 4.1 Formalization of MIAs for Diffusion Models
General Paradigm. Attacks such as Naive (Matsumoto et al., 2023), SecMI (Duan et al., 2023) and PIA (Kong et al., 2023), relying on reconstruction errors, have demonstrated effectiveness in diffusion models. These attacks share common characteristics, utilizing the model’s image prediction capability to execute attacks. They can be unified as follows: given the image to be tested $x_{i}$, calculate the distance between the image $x_{i,t}$ predicted by the model at step $t$ and the target result $x_{i,t}^{target}$ of $x_{i}$ at step $t$, then using the distance as the membership score. $x_{i,t}$ and $x_{i,t}^{target}$ are obtained in different ways depending on the attack algorithm. Then, the attacker sets a threshold $\tau$. If the score is less than $\tau$, the image is classified as member data; otherwise, it is classified as hold-out data. Formally, these attacks can be formulated as the general paradigm:
@@
4.2 Frequency Perspective of MIAs
Frequency Characteristics of Diffusion Models. Existing attacks for diffusion models mainly concentrate on reconstruction errors, yet they neglect an important aspect: analyzing models’ information processing from the frequency domain. Diffusion models’ operational mechanism features distinct frequency hierarchical properties. They first denoise low-frequency signals according to the learned distribution, then utilize specific low-frequency information as prior knowledge to process high-frequency details (Qian et al., 2024). Recovering high-frequency information effectively relies not only on the model’s learned distribution but also closely ties to the image’s inherent structure and contours during denoising. Previous research (Yang et al., 2023) has shown that diffusion models exhibit more variation and uncertainty in handling high-frequency information.
@@ 4.2 Frequency Perspective of MIAs
Frequency Effects on MIAs for Diffusion. To study how frequency domain information affects MIAs, we visualized the relationship between high-frequency content and existing attacks’ membership scores. First, we transform images to the frequency domain using the Fourier transform and divide the high- and low-frequency regions by setting the frequency domain $radius=5$ (MS-COCO (Lin et al., 2014), Flickr (Young et al., 2014)) and $radius=2$ (CIFAR-100 (Krizhevsky et al., 2009), TINY-IN (Deng et al., 2009)) as the high-low frequency boundary. Then, we calculated the percentage of high-frequency content by summing the squared frequency components in both regions. As shown in Fig. 1, visualizing the scores of attacks reveals a trend that as the high-frequency content of the images increased, so did the membership scores. This shows current attacks are biased, giving higher scores to images with more high-frequency content. Higher scores mean a lower degree of fitting, making it more likely to be classified as hold-out data. We term the phenomenon "high-frequency deficiency".
Table 1: High-frequency content statistics for failed samples. Failed samples refer to those that are misclassified, including member data identified as non-member and non-member data identified as member. In the failed samples, the high-frequency content of the member data is significantly larger than that of the hold-out data.
Method CIFAR-100 TINY-IN MS-COCO Flickr
@@
4.3 High-Frequency Filter for Enhanced MIAs
Based on the above observations, the following conclusion can be drawn: existing MIAs have not adequately accounted for the effect of high-frequency deficiency, which leads to their failure in certain specific scenarios. Intuitively, the inherent deficiency of diffusion models in handling high-frequency components leads to confusion in distinguishing between images with different high-frequency content. Further, we explore the effects of high-frequency deficiency from a theoretical perspective. We rely on the previously established definition of membership inference attack capability (Yeom et al., 2018). Definition 4.1 gives an advantage measure that characterizes how well an algorithm can distinguish between member data and hold-out data.
Definition 4.1.
The membership advantage of attack algorithm $\mathcal{A}$ is defined as:
@@ 4.3 High-Frequency Filter for Enhanced MIAs
where $\Pr[\mathcal{A}=1|m=1]$ indicates the probability that attack algorithm $\mathcal{A}$ identifies the data as a member when the data is a member. $\Pr[\mathcal{A}=1|m=0]$ indicates the probability that if the data is hold-out data, algorithm $\mathcal{A}$ will identify it as a member.
When membership scores follow a normal distribution, (Yeom et al., 2018) reinterpret the membership advantage in terms of distributional differences. Specifically, they express the advantage using the ratio between the standard deviation of the hold-out scores $\sigma_{H}$ and that of the member scores $\sigma_{M}$. And they further demonstrate that this ratio is positively correlated with the membership advantage:
@@
A larger ratio $\sigma_{H}/\sigma_{M}$ corresponds to a greater advantage for the attacker. When $\sigma_{H}<\sigma_{M}$, it means that there is no member advantage. Mathematically, the high-frequency deficiency contributes equally to a high standard deviation in the scores of both member and hold-out data, which may lead to a decrease in $\sigma_{H}/\sigma_{M}$, weakening the advantages of identifying member data and consequently interfering with the effective differentiation between member and hold-out data. This perspective is validated in Proposition 4.2.
@@
$\beta_{i,t}(r)=s\quad\text{if }r>r_{t},\;\text{otherwise }1,$ (10)
@@
Furthermore, we theoretically demonstrate in Proposition 4.2 that filtering out high-frequency components enhances the attack capability of the algorithm.
Proposition 4.2.
Assuming the attack has a membership advantage $Adv^{M}(\mathcal{A})$. Denote the original standard deviations of its membership scores in member and hold-out data as $\sigma_{M}$ and $\sigma_{H}$. And the standard deviations after removing the high-frequency components are $\sigma_{M}^{\prime}$ and $\sigma_{H}^{\prime}$. The standard deviation of membership scores in the high-frequency components is $h_{M}(h_{H})$, and the low-frequency components is $l_{M}(l_{H})$ in member(hold-out) data. Let $l_{H}-l_{M}=\Delta$, $h_{M}=k\cdot h_{H}$ with $k>0$. If $k^{2}>1+\frac{2\Delta}{h_{H}^{2}}(l_{M}+2\Delta-\sqrt{(l_{M}+2\Delta)^{2}+h_{H}^{2}})$, we have:
@@
We provide the detailed proof in Appendix C. ∎
Since $l_{M}+2\Delta-\sqrt{(l_{M}+2\Delta)^{2}+h_{H}^{2}}<0$, $\sigma_{H}^{\prime}/\sigma_{M}^{\prime}>\sigma_{H}/\sigma_{M}$ constantly hold when $k\geq 1$. The key insight of Proposition 4.2 is that when the ratio of $h_{M}/h_{H}$ exceeds a certain threshold, high-frequency contents will weaken the membership advantage, and filtering the membership scores derived from high-frequency components will amplify the membership advantage. From a theoretical perspective, it demonstrates that high-frequency components reduce the attack capability of the algorithm, while also validating the effectiveness of high-frequency filtering methods.
To further validate the practical applicability of the theoretical framework, we investigated the constraint conditions of $k$. The result shows that $k$ satisfies its constraint conditions under normal circumstances. We also provide a discussion on normality and verify the validity of the normality assumption in Proposition 4.2, while experimental results further demonstrate that our method does not rely on the assumption of normality. Detailed experiments and analysis are provided in Appendix D.2.
@@
Method STL10-U CIFAR-100 Tiny-IN
@@
Method Pokémon MS-COCO Flickr
@@
5 Experiments
5.1 Experiment Setup
Baselines. We employed the SecMI (Duan et al., 2023), PIA (Kong et al., 2023), and Naive (Matsumoto et al., 2023) as our baselines for comparison. These approaches are characterized as error-based attacks and operate effectively within a grey-box setting. We adopt the default parameters and threshold acquisition procedures as specified in the original publications of each baseline.